# Learning Constrained Optimization with Deep Augmented Lagrangian Methods

### 06 Mar 2024 v1

@@ 2 Problem Overview
Feasibility in the output space of an end-to-end trainable model $\mathcal{X}_{\bm{\theta}}=\mathcal{S}\circ\mathcal{N}_{\bm{\theta}}$ is generally guaranteed by one of two approaches. A differentiable projection $\mathcal{S}$ onto the feasible set can be combined with a prediction model $\mathcal{N}_{\bm{\theta}}$ to ensure feasibility of $\mathcal{X}_{\bm{\theta}}$ in end-to-end training [22]. However, this is generally inefficient except in limited cases where the constraints have special structure admitting a fast projection [2]. To overcome this, a notable alternative to projections is proposed in [7], which instead corrects feasibility by unrolling differentiable gradient descent steps on the constraint violations of a solution estimate during training. A distinct strategy is based on reparametrization of the model’s output space to exclude infeasible solutions. However, this has only been demonstrated in cases where the constraints ((1b), (1c)) are linear or quadratic inequalities [10, 12]. No known work has shown that feasibility to arbitrary constraints in the outputs of an end-to-end trainable model can be efficiently and reliably guaranteed.
@@
Several works propose to satisfy constraints (1b, 1c) only approximately in the outputs of a DNN model $\mathcal{N}_{\bm{\theta}}$, by training it with a loss function that incorporates penalties on the violation of each constraint [8]. At inference time, the outputs of $\mathcal{N}_{\bm{\theta}}$ are post-processed with a feasibility correction step $\mathcal{S}$. This approach allows $\mathcal{S}$ to take the form of a more lightweight model, such as a few iterations of Newton’s method on the constraint violations [15], since the outputs of $\mathcal{N}_{\bm{\theta}}$ are already nearly feasible. However, from this approach arises the problem of finding appropriate penalty weights. One line of work [9, 18, 14] interprets the penalty weights as dual variables. Updating the penalty weights by update rules inspired by dual optimization improves constraint satisfaction, but their inexactness leads to inconsistent constraint satisfaction. This often leads them to rely on hyperparameter search over initial penalty weights and update rules.
This paper proposes an altogether different approach, combining concepts from the two main approaches described above. By learning to solve the dual problem, constraint satisfaction is easily maintained in the outputs of a proxy solver throughout training. Primal solution estimates are then recovered via the stationarity condition using optimization in-the-loop. This allows the dual objective function to be maximized as a training loss, by which primal solution estimates iterate toward feasibility at convergence. Before the proposed framework is decribed in detail, the next section defines preliminary notions in Lagrangian duality.
@@ 3 Preliminaries: Lagrangian Duality
The method of constrained optimization learning proposed in this paper is based on concepts of Lagrangian Duality. In particular, the proposed method is derived by analogy to the Dual Ascent algorithm. This section provides requisite background. The following notions are defined with respect to the parametric optimization (2), for some particular instantiation of the parameters $\bm{c}$.
@@ 4.1 Training Algorithm
The Lagrangian function of (9c) is as defined as in (3). These derivative rules enable a novel training scheme aimed at realizing (4). For reasons described below, it is named Deep Dual Ascent (DDA).
@@ 4.2 Dual Ascent Interpretation
To the best of the authors’ knowledge, the use of dual function gradients (4.1) in training machine learning models is novel. On the other hand, their use is fundamental to many algorithms which solve optimization problems via their dual problems. In fact, the behavior of Algorithm 1 is best understood in light of its similarities to the classical Dual Ascent method.
@@ 6.3 Results
Learning Nonconvex Optimization

### 14 Mar 2024 v2

@@ 3 Related Work
Feasibility in the output space of an end-to-end trainable model $\mathcal{X}_{\bm{\theta}}=\mathcal{S}\circ\mathcal{N}_{\bm{\theta}}$ is generally guaranteed by one of two approaches. A differentiable projection $\mathcal{S}$ onto the feasible set can be combined with a prediction model $\mathcal{N}_{\bm{\theta}}$ to ensure feasibility of $\mathcal{X}_{\bm{\theta}}$ in end-to-end training [25]. However, this is generally inefficient except in limited cases where the constraints have special structure admitting a fast projection [2]. To overcome this, a notable alternative to projections is proposed in [7], which instead corrects feasibility by unrolling differentiable gradient descent steps on the constraint violations of a solution estimate during training. Another strategy is based on reparametrization of the model’s output space to exclude infeasible solutions. However, this has only been demonstrated in cases where the constraints ((1b), (1c)) are linear or quadratic inequalities [10, 13]. No known work has shown that feasibility to arbitrary constraints in the outputs of an end-to-end trainable model can be efficiently and reliably guaranteed.
@@
Several works propose to satisfy constraints (1b, 1c) only approximately in the outputs of a DNN model $\mathcal{N}_{\bm{\theta}}$, by training it with a loss function that incorporates penalties on the violation of each constraint [8]. At inference time, the outputs of $\mathcal{N}_{\bm{\theta}}$ are post-processed with a feasibility correction step $\mathcal{S}$. This approach allows $\mathcal{S}$ to take the form of a more lightweight model, such as a few iterations of Newton’s method on the constraint violations [16], since the outputs of $\mathcal{N}_{\bm{\theta}}$ are already nearly feasible. However, from this approach arises the problem of finding appropriate penalty weights. One line of work [9, 19, 15] interprets the penalty weights as dual variables. Updating the penalty weights by update rules inspired by dual optimization improves feasibility, but their inexactness leads to inconsistent constraint satisfaction. This often leads them to rely on hyperparameter search over initial values and update steps for the penalty weights.
Learning Dual Solutions
Recently, another related direction has considered learning dual solutions to parametric optimization problems. The first such known work [20] learns dual solutions to a convex conic programming model of optimal power flow. Two further works [12] and [23] have proposed to learn dual solutions to general linear programs and convex conic programs respectively, by incorporating a dual interior point method an a fast projection onto the dual cone, respectively. In these works, the learned model employs feasibility dual predictions as fast estimates on lower bounds of the primal objective function, as certificates of optimality.
@@ 4 Preliminaries: Lagrangian Duality
The method for learning constrained optimization proposed in this paper is based on concepts of Lagrangian Duality. This section provides requisite background. The following notions are defined with respect to the parametric optimization (2), for some particular instantiation of the parameters $\bm{c}$.
@@ 5.1 Training Algorithm
The Lagrangian function of (9c) is as defined as in (3). These derivative rules enable a training scheme aimed at realizing (5) directly. For reasons described below, it is named Deep Dual Ascent (DDA).
@@ 5.2 Dual Ascent Interpretation
The use of dual function gradients (5.1) is fundamental to many algorithms which solve optimization problems via their dual problems. In fact, the behavior of Algorithm 1 is best understood in light of its similarities to the classical Dual Ascent method.
@@ 7.3 Results
Learning Nonconvex Optimization.
@@ 8 Conclusions
Acknowledgments
This research is partially supported by NSF grants 2334936, 2334448, and NSF CAREER Award 2401285. The views and conclusions of this work are those of the authors only.